# AIGC for Industrial Time Series: From Deep Generative Models to Large Generative Models

### v1

@@ VII-A Generative Models for Low-quality and Mixed-type Data
To address data quality issues, the collected data can be pre-processed before being input into the model. In addition, more robust generative models can be developed to effectively deal with quality issues of noise and outliers in the industry. For mixed data types, the generation of mixed data can be facilitated by co-representation. For instance, Lee et al.[137] proposed a co-evolving contrastive generation framework to separately model continuous and discrete variables of tabular data. During training, the two diffusion models evolve together by exchanging conditions with each other. In addition, a general framework can be developed to include the generation, estimation, and prediction of mixed-type time series models.
VII-B Generative Models with Long Temporal Expressive Power
@@
Future research could focus on exploring more advanced sequence modeling methods to handle long-term temporal dependencies more effectively. Advanced recurrent neural network (RNN) structure[138] is one such avenue. In addition, the attention mechanism[121] and state space models[139] are also effective methods. By capturing long-term temporal dependencies, the model can generate time series data more consistent with temporal correlations.
VII-C Interpretable/Reliable/Credible Generative Models
@@
VII-D Industrial Multimodal Generative Models
@@
Future research directions should focus on the development of generative models applicable to multimodal data. This may involve an in-depth study of representation learning methods for multimodal data to better capture the correlations among different sensors, which will help the models to more fully simulate the behavior and state of industrial systems. In addition, by joining modalities from different data not only can performance be enhanced but possible tasks can be explored. For instance, AnomalyGPT[140] integrates the information of Industrial Anomaly Detection (IAD) task to the Large Vision-Language Models (LVLMs) to support the multi-dialogues. Moreover, it doesn’t need manual threshold adjustment, which can directly present the anomalies. Stable diffusion[141] generates stunning images from text prompts alone, dramatically reducing the amount of manual labor required. Multimodal time series generation can be performed by combining text and signals.
VII-E Industrial Generalized Generative Models
@@
VII-F Generative Models for Smart Manufacturing
Generative modeling can generate a variety of industrial design prototypes to help engineers explore diverse device concepts and accelerate product design and innovation, including code generation[144, 145, 146], production process generation[147, 148, 149], and so on. The advantages of DGMs include: (1) Generative design models can produce diverse designs and stimulate the creativity of designers. (2) Automated code generation can reduce the cost of generation and design, especially in mass production. For instance, the paper[146] proposed a generative AI methodology that automatically translates nuclear power plant safety operation flowcharts into executable code to improve the efficiency and reduce the cognitive burden of operators using flowcharts.
Currently, LGMs and DGMs have received much attention due to their ability to understand, generate, and create data. For this reason, generative models have made significant progress and show great potential in industrial time series. In conclusion, this review provides a detailed overview of the generative models in industrial. First, we analyze the definition, characteristics, and data scarcity challenges within industrial time series generation. Then, we propose a DGM-based AIGC framework for industrial data generation and present a multi-perspective categorization of industrial DGMs. We further discuss the application of DGMs in industrial time series. In addition, we explore the construction of LGMs in the industrial domain. Finally, we conclude with challenges and future directions to enable the development of generative models in industrial applications.

### v2

@@ IX-A Generative Models for Low-quality and Mixed-type Data
To address data quality issues, the collected data can be pre-processed before being input into the model. In addition, more robust generative models can be developed to effectively deal with quality issues of noise and outliers in the industry. For mixed data types, the generation of mixed data can be facilitated by co-representation. For instance, Lee et al.[135] proposed a co-evolving contrastive generation framework to separately model continuous and discrete variables of tabular data. In addition, a general framework can be developed to include the generation, estimation, and prediction of mixed-type time series models.
IX-B Generative Models with Long Temporal Expressive Power
@@
Future research could focus on exploring more advanced sequence modeling methods to handle long-term temporal dependencies more effectively. Advanced recurrent neural network (RNN) structure is one such avenue. In addition, the attention mechanism[118] and state space models[136] are also effective methods.
IX-C Interpretable/Reliable/Credible Generative Models
@@
IX-D Industrial Multimodal Generative Models
@@
Future research directions should focus on the development of generative models applicable to multimodal data. This may involve an in-depth study of representation learning methods for multimodal data to better capture the correlations among different sensors, which will help the models to more fully simulate the behavior and state of industrial systems. In addition, by joining modalities from different data not only can performance be enhanced but possible tasks can be explored. For instance, AnomalyGPT[137] integrates the information of Industrial Anomaly Detection (IAD) task to the Large Vision-Language Models (LVLMs) to support the multi-dialogues. Moreover, it doesn’t need manual threshold adjustment, which can directly present the anomalies. Multimodal time series generation can be performed by combining text and signals.
IX-E Industrial Generalized Generative Models
@@
IX-F Generative Models for Smart Manufacturing
Generative modeling can generate a variety of industrial design prototypes to help engineers explore diverse device concepts and accelerate product design and innovation, including code generation[140], production process generation, and so on. The advantages of DGMs include: (1) Generative design models can produce diverse designs and stimulate the creativity of designers. (2) Automated code generation can reduce the cost of generation and design, especially in mass production. For instance, the paper[141] proposed a generative AI methodology that automatically translates nuclear power plant safety operation flowcharts into executable code to improve the efficiency and reduce the cognitive burden of operators using flowcharts.
IX-G Ethical and Privacy Consideration in Generative Models
The deployment of generative models may introduce critical ethical and privacy challenges. The first issue is algorithmic bias. When generative models are trained on historical industrial data, they can inadvertently inherit and even amplify existing biases. To address this, techniques such as fairness-aware training[142] and adversarial debiasing[143] have proven effective in reducing correlations between generated outputs and sensitive attributes, ensuring that the model does not perpetuate historical inequities. These approaches help recalibrate the model’s learning process, promoting a more balanced and just representation of the underlying industrial data.
The second issue is data privacy and security. Industrial datasets may contain sensitive, proprietary information, and the misuse or leakage of such data could have severe consequences. Integrating differential privacy into the training process can ensure that individual data points remain unidentifiable, even in the synthetic outputs generated by the model. Furthermore, adopting federated learning frameworks[144] allows multiple stakeholders to collaboratively train models without the need to exchange raw data, thereby preserving data locality and enhancing security.
IX-H Conclusion
Currently, LGMs and DGMs have received much attention due to their ability to understand, generate, and create data. For this reason, generative models have made significant progress and show great potential in industrial time series. In conclusion, this review provides a detailed overview of the generative models in industrial. First, we analyze the definition, characteristics, and data scarcity challenges within industrial time series generation. Then, we propose a DGM-based AIGC framework for industrial data generation and present a multi-perspective categorization of industrial DGMs. We further discuss the application of DGMs in industrial time series. Additionally, we establish a comprehensive evaluation benchmark for generative models, systematically assessing fidelity, diversity, and utility. In addition, we explore the construction of LGMs in the industrial domain. Finally, we conclude with challenges and future directions to enable the development of generative models in industrial applications.